OpenStudy (anonymous):

find the domain of the function. h(x)=10/x^2-2x

OpenStudy (anonymous):

Is your \[h(x)=\frac{10}{x^2-2x}\]function

OpenStudy (anonymous):

yes that equation

OpenStudy (anonymous):

It can be rewritten as\[h(x)=\frac{10}{x(x-2)}\] The function's domain includes all numbers where the denominator is not 0. In this case, the denominator is 0 when x=0 or 2. So the domain is \[(-\infty, 0), (0,2), (2, \infty)\]

OpenStudy (anonymous):

Then the domain is all real numbers except 0 and 2. alexray has shown you how to factor the denominator and has given the domain in interval notation.